OpenStudy (anonymous):

how come pi/4 equals sqrt2/2?

OpenStudy (usukidoll):

rationalize the denominator

OpenStudy (anonymous):

unit circle

OpenStudy (usukidoll):

actually only sinx and cosx has sqrt(2) / 2 tan 45 is 1

OpenStudy (callisto):

Maybe you mean cos(pi/4) or sin(pi/4) instead of pi/4 ?

OpenStudy (anonymous):

yes cos pi/4

OpenStudy (anonymous):

why does it equal sqrt2/2?

OpenStudy (anonymous):

look at the unit circle, pi/4 = sqrt (2) / 2 for the cos value the coordinates for the polar coordinate system is (cos,sin).

OpenStudy (usukidoll):

draw the right triangle of 45 degrees.

OpenStudy (anonymous):

http://image.tutorvista.com/cms/images/38/unit-circle.JPG You can see here, that at 45 degrees or pi /4 . The value for cosine is sqrt 2 / 2

OpenStudy (usukidoll):

if only draw WOULD WORK ON HERE THAT WOULD BE GREATTTTTTTTTTTTTT

OpenStudy (callisto):

What is the length of the hypotenuse?

OpenStudy (usukidoll):

missing sqrt(2)

OpenStudy (callisto):

And you can then find the value of cos(pi/4)

OpenStudy (callisto):

@UsukiDoll I was asking him to find the missing side.

OpenStudy (usukidoll):

uh oh! sorry X<X

OpenStudy (anonymous):

the problem says use the fact that trig functions are periodic to find the exact value of cos 17pi/4

OpenStudy (usukidoll):

ok first multiply by 180 / pi

OpenStudy (anonymous):

I can't draw, damn it! I need drawing to freaking do trig.

OpenStudy (usukidoll):

same here ! ! !

OpenStudy (anonymous):

I can't draw the triangles

OpenStudy (usukidoll):

IT'S TWIDDLA TIME! :D

OpenStudy (usukidoll):

http://www.twiddla.com/1470501

OpenStudy (usukidoll):

there now we can draw!

OpenStudy (agent0smith):

Subtract multiples of 2pi from 17pi/4... since 2pi is a full revolution 17pi/4 - 16pi/4 =

OpenStudy (agent0smith):

16pi/4 = 4pi, a multiple of 2pi

OpenStudy (anonymous):

hypotenuse = 1 first leg = x second leg = y We know x^2 + y^2 = 1^2 We also know that x = y x^2 + x^2 = 1^2 2 * x^2 = 1 x^2 = 1/2 x = 1/sqrt(2)

OpenStudy (anonymous):

I know that I have to find the biggest multiple of 2 smaller than 17/4 then find the number to complete which would be pi/4 I just don't know why it equals sqrt2/2

OpenStudy (anonymous):

When you rationalize the denominator: 1/sqrt(2) * sqrt(2) / sqrt(2) = sqrt(2) / 2

OpenStudy (anonymous):

is my suggestion being completely ignored? -_- after you do what agent does, you look at the unit circle... polar coordinates are in (cos,sin) form. pi/4 on the unit circle is (sqrt(2)/sqrt(2)/2) http://image.tutorvista.com/cms/images/38/unit-circle.JPG

OpenStudy (usukidoll):

It's much simpler than that D:

OpenStudy (anonymous):

(sqrt(2)/2,sqrt(2)/2) *

OpenStudy (anonymous):

Looking at the unit circle is much simpler than doing what you are doing..

OpenStudy (usukidoll):

we agree to disagree... I don't want to get more warnings.

OpenStudy (anonymous):

The unit circle is blasphemous in this question.

OpenStudy (usukidoll):

OOOOOOOOOOOOO SNAP!

OpenStudy (anonymous):

i'm already banned from community, lul

OpenStudy (campbell_st):

isn't cos someone said so....

OpenStudy (anonymous):

ok so 17//4=4.25 so the biggest multiple of 2pi smaller than 4.25 is 4 so 0.25 is 1/4 but Im looking for the formula to get from pi/4 to sqrt2/2

OpenStudy (anonymous):

There is no simple formula for it.

OpenStudy (anonymous):

There are ways to calculate it, for sure, but you can't just put it into some simple formula.

OpenStudy (anonymous):

thanks everyone